OpenStudy (anonymous):

adam earned 1 1/2 times his regular rate of $81/2 for 3 hours. how much did he earn

OpenStudy (anonymous):

8.5/2 = 4.25 or 41/4 81/2 + 41/4 = 12.75 OR 12 3/4 12.75 * 3 = 38.25 or 381/4